OpenStudy (anonymous):

Owen scored eighty-eight points on a one hundred-point test and twenty-five on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

Just divide both of them and see what you get: \[\frac{88}{100} = \frac{22}{25} = 0.88\] \[\frac{25}{30} = \frac{5}{6} = 0.83\] Which one is greater ??

OpenStudy (anonymous):

0.88

OpenStudy (anonymous):

Well Done..